OpenStudy (anonymous):

Write y =1/6 x + 5 in standard form using integers.

OpenStudy (experimentx):

it is in standard form y = mx+c i guess you mean, 6y = x + 30

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (experimentx):

it depends on your y-intercept ... if the line is passing through origin, i.e. y=0 when x=0 then it would be -2x8/-6 = 8/3

OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (phi):

Here are the definitions for the various equations of a line http://en.wikipedia.org/wiki/Equation_of_a_line#Forms_for_2D_linear_equations standard form is Ax+By=C where A is positive.

OpenStudy (phi):

Direct variation means can you find a number K so x = K y if you have 5x= -3y can you get x by itself on the left?

OpenStudy (experimentx):

i guess so

OpenStudy (phi):

try it

OpenStudy (phi):

divide both sides by 5

OpenStudy (anonymous):

A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5 these are some of the choices

OpenStudy (anonymous):

???

OpenStudy (phi):

Direct variation means can you find a number K so x = K y if you have 5x= -3y can you get x by itself on the left? divide both sides by 5